# PRoTECT: Parallelized ConstRuction of SafeTy BarriEr Certificates for Nonlinear Polynomial SysTems

### v1

@@ Abstract.
We develop an open-source software tool, called PRoTECT, for the parallelized construction of safety barrier certificates (BCs) for nonlinear polynomial systems. This tool employs sum-of-squares (SOS) optimization programs to systematically search for polynomial-type BCs, while aiming to verify safety properties over four classes of dynamical systems: (i) discrete-time stochastic systems, (ii) discrete-time deterministic systems, (iii) continuous-time stochastic systems, and (iv) continuous-time deterministic systems. PRoTECT is implemented in Python as an application programming interface (API), offering users the flexibility to interact either through its user-friendly graphic user interface (GUI) or via function calls from other Python programs. PRoTECT leverages parallelism across different barrier degrees to efficiently search for a feasible BC.
@@ 1.0.1. Motivation for PRoTECT
Barrier certificates, initially introduced in [25, 26], have gained significant recognition for the formal safety analysis of dynamical systems over the past 20 years. These approaches can verify systems with polynomial dynamics, enabling the design of polynomial BCs using sum-of-squares techniques, e.g., via SOSTOOLS [27]. Additionally, some alternative approaches explore verifying nonpolynomial systems, such as those discovered through counter-example guided inductive synthesis (CEGIS), leveraging satisfiability modulo theories (SMT) solvers including Z3 [9], dReal [11], or MathSat [8]. Other new techniques encompass neural barrier functions [35, 20] and genetic programs [32]. While we primarily focus on utilizing BCs for safety specifications in PRoTECT, they also hold significant value for addressing other specifications, including reachability, reach-while-avoid, and other temporal logic specifications [5, 19]. A comprehensive overview of barrier certificates in both deterministic and stochastic settings can be found in the line of work [4, 33, 21, 22, 16, 12].
@@
To the best of our knowledge, there exists only one other tool dedicated to the construction of barrier certificates, FOSSIL [1]. In the latest release [10], FOSSIL has expanded its capabilities to include finding barrier certificates for discrete- and continuous-time deterministic systems using the CEGIS approach, while facilitating verification and control synthesis for specifications including safety, reachability, and reach-while-avoid. However, FOSSIL lacks support for stochastic systems, whereas in PRoTECT, we provide support for both discrete- and continuous-time stochastic systems. Additionally, while FOSSIL can handle nonpolynomial BCs, it does not guarantee termination due to its use of CEGIS approach. In contrast, PRoTECT utilizes sum-of-squares optimization techniques, enabling the exploitation of parallelism to concurrently search for multiple candidate BCs with different degrees, aiming to construct them effectively.
There have also been some tools that concentrate on finite abstraction methods, an approach capable of providing safety guarantees but hampered by the curse of dimensionality, stemming from the necessity to grid the state/input space. Typically, these tools focus on only one or two cases of dynamical systems (e.g., SCOTS [30] or AMYTISS [17]); however, in PRoTECT we consider four classes of dynamical systems.
@@
    We propose the first tool, employing SOS optimization techniques, that verifies the safe behavior of four classes of dynamical systems: (i) discrete-time stochastic systems (dt-SS), (ii) discrete-time deterministic systems (dt-DS), (iii) continuous-time stochastic systems (ct-SS), and (iv) continuous-time deterministic systems (ct-DS).
@@
    In stochastic settings, PRoTECT automatically optimizes the safety confidence level while designing different decision variables, including level sets of BCs.
@@ 3. Discrete-Time Stochastic Systems
$\mathcal{B}(x)\leq\gamma,\quad$ $\forall x\in X_{\mathcal{I}},$ (2)
$\mathcal{B}(x)\geq\lambda,\quad$ $\forall x\in X_{\mathcal{U}},$ (3)
$\mathbb{E}\Big{[}\mathcal{B}(f(x,\varsigma))~{}\big{|}~{}x\Big{]}\leq\mathcal{B}(x)+c,\quad$ $\forall x\in X,$ (4)
@@
We now leverage the BC in Definition 2 and quantify a lower bound confidence over the safety of dt-SS [15, 26].
@@ 3.1. PRoTECT Implementation for dt-SS
3.1.1. Graphic User Interface (GUI)
@@
3.1.2. Application Programming Interface (API).
In general, the backend of PRoTECT behaves as an API, with functions that can be called and used in any python program. We provide some generic configuration files in /ex/benchmarks-stochastic and /ex/benchmarks-deterministic, which demonstrate how to use the functions in a standard python program. The user is expected to provide the following required parameters: dimension of the state set $X\subseteq\mathbb{R}^{n}$ (blue-2), indicated by dim, and the degree of the barrier certificate (blue-3), denoted by b_degree. The lower and upper bounds of the initial region $X_{\mathcal{I}}$ , labeled as L_initial and U_initial; lower and upper bounds of the unsafe region $X_{\mathcal{U}}$, referred to as L_unsafe and U_unsafe; lower and upper bounds for the state set $X$, denoted as L_space and U_space; where the value of each dimension is separated with a comma (blue-4). Due to possible scenarios with multiple unsafe regions, the unsafe region is passed to the functions as a numpy array of numpy arrays describing each individual unsafe region. The transition map $f$, represented by f, written as a SymPy expression [^1] https://docs.sympy.org/latest/tutorials/intro-tutorial/basic_operations.html for each dimension using states x1,x2,... and noise parameters varsigma1,varsigma2,... (blue-5). The time horizon $\mathcal{T}$, noted as t (blue-7). The distribution of the noise, NoiseType, can be specified as either "normal", "exponential", or "uniform" (blue-6).
@@ 4. Discrete-Time Deterministic Systems
$\mathcal{B}(f(x))\leq\mathcal{B}(x),\quad\forall x\in X.$ (9)
@@ 5. Continuous-Time Stochastic System
$\mathcal{L}\mathcal{B}(x)\leq c,\quad\forall x\in X,$ (11)
@@
One can apply Lemma 1 to offer the confidence level $\phi$ in (5) using the constructed level sets $\gamma,\lambda$ and the constant $c$ in (12).
@@ 6. Continuous-Time Deterministic Systems
Continuous-time deterministic systems (ct-DS), as the last class of models, can be described by $\Sigma_{c}:\dot{x}(t)=f(x(t)),\quad t\in\mathbb{R}^{+}_{0}$ where $f:X\rightarrow X$ is the vector field. A differentiable BC $\mathcal{B}:X\rightarrow\mathbb{R}$ for ct-DS is defined akin to Definition 2, maintaining conditions (2)-(3), with an adjustment to condition (4) with $c=0$ as
$\text{L}_{f}\mathcal{B}(x)\leq 0,\quad\forall x\in X,$ (13)
@@ 7. Benchmarking and Case Studies
Table 1 presents a comparison of the efficiency in finding barrier certificates for deterministic systems, dt-DS and ct-DS, using PRoTECT against the tool FOSSIL [1]. Unless referenced, the case studies are taken from FOSSIL’s own benchmarks. Descriptions of all case studies, along with their corresponding simulation results, are provided in the appendix. All examples in this table, including the FOSSIL configuration files, can be located in the folder /ex/benchmarks-deterministic. Notably, PRoTECT demonstrates higher efficiency for case studies up to four dimensions, while FOSSIL excels for those with six and eight dimensions. This discrepancy can be attributed to the increase in Lagrangian multipliers in PRoTECT, which also necessitates designing more decision variables when searching for higher-degree BCs. Table 2 employs PRoTECT for stochastic benchmarks. We remind the reader that, unlike the deterministic setting, PRoTECT awaits barriers of all degrees to terminate, returning the one with the highest confidence. Consequently, stochastic case studies typically require notably longer completion times compared to deterministic ones, but at the gain of offering the highest confidence.
Table 1. Efficiency comparison of BC construction for deterministic systems: PRoTECT vs. FOSSIL. Models marked with “” were adjusted to switch domains from spheres to hypercubes. All case studies were run on the same desktop computer (Intel i9-12900).
@@
$n$ system $\gamma$ $\lambda$ serial (sec) parallel (sec) (sec)
$1$D system 1 dt-DS 3.90 4.04 0.09 1.16 0.20
DC Motor [3] 2 dt-DS 1.19 1.20 0.20 1.31 0.22
barr2roomDT 2 dt-DS 18.0 19.5 0.17 1.32 0.44
Jet Engine [14] 2 ct-DS 1.41 1.53 0.18 0.22 0.28
hi-ord4 4 ct-DS 14.5 15.0 1.30 1.42 27.5
hi-ord6 6 ct-DS 32.5 33.9 17.0 17.6 11.6
hi-ord8 8 ct-DS 45.7 57.2 172 172 21.3

### v2

@@ Abstract.
We develop an open-source software tool, called PRoTECT, for the parallelized construction of safety barrier certificates (BCs) for nonlinear polynomial systems. This tool employs sum-of-squares (SOS) optimization programs to systematically search for polynomial-type BCs, while aiming to verify safety properties over four classes of dynamical systems: (i) discrete-time stochastic systems, (ii) discrete-time deterministic systems, (iii) continuous-time stochastic systems, and (iv) continuous-time deterministic systems. In particular, PRoTECT is the first software tool that offers stochastic barrier certificates. PRoTECT is implemented in Python as an application programming interface (API), offering users the flexibility to interact either through its user-friendly graphic user interface (GUI) or via function calls from other Python programs. PRoTECT leverages parallelism across different barrier degrees to efficiently search for a feasible BC.
@@ 1. Introduction
Barrier certificates, initially introduced in [25, 26], have gained significant recognition for the formal safety analysis of dynamical systems over the past 20 years. These approaches can verify systems with polynomial dynamics, enabling the design of polynomial BCs using sum-of-squares techniques, e.g., via SOSTOOLS [27]. Additionally, some alternative approaches explore verifying nonpolynomial systems, such as those discovered through counter-example guided inductive synthesis (CEGIS), leveraging satisfiability modulo theories (SMT) solvers including Z3 [10], dReal [12], or MathSat [9]. Other new techniques encompass neural barrier functions [34, 21] and genetic programs [31]. A comprehensive overview of barrier certificates can be found in [4, 32]. While we primarily focus on utilizing BCs for safety specifications in PRoTECT, they also hold significant value for addressing other specifications, including reachability, reach-while-avoid, and other temporal logic specifications [5, 20].
@@
FOSSIL [1] is a relevant tool for the model-based construction of barrier certificates. In the latest release [11], FOSSIL has expanded its capabilities to include finding barrier certificates for discrete- and continuous-time deterministic systems using the CEGIS approach, while facilitating verification and control synthesis for specifications including safety, reachability, and reach-while-avoid. However, FOSSIL lacks support for stochastic systems, whereas in PRoTECT, we provide support for both discrete- and continuous-time stochastic systems. Additionally, while FOSSIL can handle nonpolynomial BCs, it does not guarantee termination due to its use of CEGIS approach. In contrast, PRoTECT utilizes sum-of-squares optimization techniques, enabling the exploitation of parallelism to concurrently search for multiple candidate BCs with different degrees, aiming to construct them effectively. Recently, two new tools for constructing barrier certificates have been introduced. TRUST [13] is a data-driven tool that generates barrier certificates for deterministic systems with unknown polynomial dynamics, using only a single trajectory of collected data. In addition, CBFKIT [8] is a toolbox designed for safe robotic planning. It supports both deterministic and stochastic continuous-time dynamics but requires the user to provide a barrier function a priori, which it then verifies for correctness—unlike PRoTECT, which automatically synthesizes a barrier certificate to meet the required conditions.
@@
    We propose the first tool, employing SOS optimization techniques, that verifies the safe behavior of four classes of dynamical systems: (i) discrete-time stochastic systems (dt-SS), (ii) discrete-time deterministic systems (dt-DS), (iii) continuous-time stochastic systems (ct-SS), and (iv) continuous-time deterministic systems (ct-DS). In particular, PRoTECT is the first software tool that offers stochastic barrier certificates.
@@ 3. Discrete-Time Stochastic Systems
$\mathcal{B}(x)\leq\gamma,\quad\quad$ $\forall x\in X_{\mathcal{I}},$ (2)
$\mathcal{B}(x)\geq\lambda,\quad\quad$ $\forall x\in X_{\mathcal{U}},$ (3)
$\mathbb{E}\Big{[}\mathcal{B}(f(x,\varsigma))~{}\big{|}~{}x\Big{]}\leq\mathcal{B}(x)+c,\quad\quad$ $\forall x\in X,$ (4)
@@
We now leverage the BC in Definition 2 and quantify a lower bound confidence over the safety of dt-SS [17, 26]. This lemma, commonly found in the literature (e.g. [30]), provides the safety confidence for stochastic systems. The same confidence formula applies to continuous-time stochastic systems in Section 5.
@@
This lemma is well-established in the literature and extends naturally from works such as [24, 26, 19, 14].
@@ 3.1. PRoTECT Implementation for dt-SS
3.1.1. Graphic User Interface (GUI).
@@
3.1.2. Application Programming Interface (API)
In general, the backend of PRoTECT behaves as an API, with functions that can be called and used in any python program. In the project folders /ex/benchmarks-stochastic and /ex/benchmarks-deterministic, we provide some generic configuration files which demonstrate how to use the functions in a standard python program. The user is expected to provide the following required parameters: dimension of the state set $X\subseteq\mathbb{R}^{n}$ (blue-2), indicated by dim, and the degree of the barrier certificate (blue-3), denoted by b_degree. The lower and upper bounds of the initial region $X_{\mathcal{I}}$ , labeled as L_initial and U_initial; lower and upper bounds of the unsafe region $X_{\mathcal{U}}$, referred to as L_unsafe and U_unsafe; lower and upper bounds for the state set $X$, denoted as L_space and U_space; where the value of each dimension is separated with a comma (blue-4). Due to possible scenarios with multiple unsafe regions, the unsafe region is passed to the functions as a numpy array of numpy arrays describing each individual unsafe region. The transition map $f$, represented by f, written as a SymPy expression [^1] https://docs.sympy.org/latest/tutorials/intro-tutorial/basic_operations.html for each dimension using states x1,x2,... and noise parameters varsigma1,varsigma2,... (blue-5). The time horizon $\mathcal{T}$, noted as t (blue-7). The distribution of the noise, NoiseType, can be specified as either "normal", "exponential", or "uniform" (blue-6).
@@ 4. Discrete-Time Deterministic Systems
$\mathcal{B}(f(x))\leq\mathcal{B}(x),\quad\quad\forall x\in X.$ (9)
@@ 5. Continuous-Time Stochastic System
$\mathcal{L}\mathcal{B}(x)\leq c,\quad\quad\forall x\in X,$ (11)
@@
For ct-SS, one can adapt Lemma 1 to offer the same formulation of confidence level $\phi$ in (5), using the constructed level sets $\gamma,\lambda$, and the constant $c$ in (12).
@@ 6. Continuous-Time Deterministic Systems
Continuous-time deterministic systems (ct-DS), as the last class of models, can be described by $\Sigma_{c}:\dot{x}=f(x(t)),\quad t\in\mathbb{R}^{+}_{0}$ where $f:X\rightarrow X$ is the vector field. A differentiable BC $\mathcal{B}:X\rightarrow\mathbb{R}$ for ct-DS is defined akin to Definition 2, maintaining conditions (2)-(3), with an adjustment to condition (4) with $c=0$ as
$\text{L}_{f}\mathcal{B}(x)\leq 0,\quad\quad\forall x\in X,$ (13)
@@ 7. Benchmarking and Case Studies
We highlight specific case studies conducted using PRoTECT, with their performance detailed in the following subsection. Additional case studies are provided in the appendix.
7.1. Discrete-time Stochastic Systems (dt-SS)
7.1.1. Van der Pol Oscillator.
Figure 3. $2$D Van der Pol oscillator system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(0,4)$ for a time horizon $\mathcal{T}=5$, with time increments of $0.01$. The uniform noise has bounds [-0.1,0.1] in both dimensions.
7.2. Discrete-time Deterministic Systems (dt-DS)
7.2.1. DC Motor.
Figure 4. $2$D DC Motor system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(0.3,0.2)$ for a time horizon of 100 steps.
We consider the following DC motor case study [3]
7.3. Continuous-time Stochastic Systems (ct-SS)
7.3.1. Nonlinear System.
Figure 5. $2$D nonlinear system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(0,0)$ for a time horizon $\mathcal{T}=10$ and time increments $0.01$.
We consider the following nonlinear system [25]
$\Sigma_{c}^{\delta}:\begin{cases}\mathbf{d}{x}_{1}(t)=x_{2}(t)\mathbf{d}t,\\ \mathbf{d}{x}_{2}(t)=(-x_{1}(t)-x_{2}(t)-0.5x_{1}^{3}(t))\mathbf{d}t+\delta\mathbf{d}\mathbb{W}_{t},\par\end{cases}$
7.4. Continuous-time Deterministic Systems (ct-DS)
7.5. Jet Engine Model.
Figure 6. $2$D Jet Engine system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(0.2,0.2)$ for a time horizon $\mathcal{T}=20$.
We consider the nonlinear Moore-Greitzer jet engine model [16]:
7.6. Benchmarking
Table 1. Efficiency comparison of BC construction for deterministic systems: PRoTECT vs. FOSSIL. All case studies were run on the same desktop computer (Intel i9-12900), and PRoTECT cases were run exclusively for the degree $2$ barrier.
One Shot Deterministic Systems
@@ 7.6. Benchmarking
$n$ system b_degree $\gamma$ $\lambda$ (sec) (sec)
$1$D system 1 dt-DS 2 3.90 4.04 0.09 0.20
DC Motor [3] 2 dt-DS 2 1.19 1.20 0.20 0.22
barr2roomDT 2 dt-DS 2 18.0 19.5 0.17 0.44
Jet Engine [16] 2 ct-DS 2 1.41 1.53 0.18 0.28
hi-ord4 4 ct-DS 2 14.5 15.0 1.30 27.5
hi-ord6-1 6 ct-DS 2 32.5 33.9 17.0 11.6
hi-ord6-2 6 ct-DS 2 23.73 23.74 18.9 652
hi-ord8-1 8 ct-DS 2 45.7 57.2 172 21.3
hi-ord8-2 8 ct-DS 2 5.78 8.77 175 569
Table 2. Efficiency evaluation in BC construction for stochastic systems via PRoTECT. We present case studies for the barrier degrees 4, optimizing for the barrier with the highest confidence. Parameters $\gamma$ and $c$ are designed via SOS optimization and $\lambda$ is fixed a-priori. VDP denotes the stochastic Van der Pol oscillator. All experiments were conducted on a desktop computer (Intel i9-12900).
One Shot Stochastic Systems
$n$ system $\mathcal{T}$ b_degree $\gamma$ $\lambda$ $c$ $\phi$ time (sec)
RoomTemp [22] 1 ct-SS 5 4 $4.8e^{-5}$ 10 $9.6e^{-6}$ 0.99 0.16
RoomTemp [22] 1 dt-SS 5 4 $4.4e^{-5}$ 10 $8.9e^{-6}$ 0.99 0.25
VDP [2] 2 dt-SS 5 4 N/A 1000 N/A N/A 1.65
ex_lin1 [25] 2 ct-SS 5 4 1.39 10 0.26 0.73 0.61
ex_nonlin1 [25] 2 ct-SS 5 4 3.34 10 0.53 0.40 0.61
TwoTanks [28] 2 dt-SS 5 4 $1.0e^{-6}$ 10 $3.4e^{-8}$ 0.99 1.25
RoomTemp [30] 3 dt-SS 3 4 $1.1e^{-8}$ 10 $7.5e^{-9}$ 0.99 29.1
hi-ord4 [1] 4 ct-SS 3 4 0.02 10 0.02 0.99 68.1
Table 3. Efficiency evaluation in BC construction for stochastic systems via PRoTECT. We present case studies across three barrier degrees (2, 4, and 6), returning the barrier with the highest confidence. Parameter $\lambda$ is fixed to a certain value and then $\gamma$ and $c$ are designed via SOS optimization. VDP denotes the stochastic Van der Pol oscillator. All experiments were conducted on a desktop computer (Intel i9-12900).
Parallel Stochastic Systems
$n$ system $\mathcal{T}$ b_degree $\gamma$ $\lambda$ $c$ $\phi$ time (sec)
RoomTemp [22] 1 ct-SS 5 6 $1.1e^{-6}$ 10 $2.3e^{-7}$ 0.99 0.33
RoomTemp [22] 1 dt-SS 5 6 $4.4e^{-7}$ 10 $1.0e^{-7}$ 0.99 0.53
VDP [2] 2 dt-SS 5 6 97.5 1000 3.53 0.88 14.3
ex_lin1 [25] 2 ct-SS 5 6 0.34 10 0.04 0.95 1.73
ex_nonlin1 [25] 2 ct-SS 5 6 1.84 10 0.2 0.72 1.81
TwoTanks [28] 2 dt-SS 5 4 $1.0e^{-6}$ 10 $3.4e^{-8}$ 0.99 5.14
RoomTemp [30] 3 dt-SS 3 4 $1.1e^{-8}$ 10 $7.5e^{-9}$ 0.99 1501
hi-ord4 [1] 4 ct-SS 3 6 $1.8e^{-3}$ 10 $1.2e^{-3}$ 0.99 1308
Table 1 compares the efficiency of PRoTECT and FOSSIL [1] in finding barrier certificates for deterministic systems, both dt-DS and ct-DS. PRoTECT was operated in a one-shot mode, searching for a barrier certificate without parallelism. Unless referenced, the case studies are taken from FOSSIL’s own benchmarks. All examples in this table, including the FOSSIL configuration files, can be located in the folder /ex/benchmarks-deterministic. Notably, PRoTECT demonstrates always better time efficiency for case studies up to four dimensions, while FOSSIL may perform better in cases with six and eight dimensions. This can be attributed to the increase in Lagrangian multipliers in PRoTECT, which also necessitates designing more decision variables when searching for higher-degree BCs. However, as demonstrated in the hi-ord6-2 and hi-ord8-2 cases, FOSSIL performs significantly worse than PRoTECT, which exhibits superior time efficiency in solving these problems. This is primarily because FOSSIL relies on a neural network approach to solve the problem, and the tool’s runtime can vary depending on the network’s structure and the corresponding parameters involved, as well as the verification process using SMT solvers. We re-highlight that FOSSIL does not guarantee convergence to a solution.
Table 2 and Table 3 employ PRoTECT for stochastic benchmarks. In the ‘One Shot’ setting with a fixed barrier degree of $4$, as shown in Table 2, PRoTECT optimizes all parameters, $\gamma$ and $c$, during the SOS formulation, while $\lambda$ is fixed a priori. Notably, the VDP example cannot find a barrier certificate with degree $4$. We do not provide any comparison with FOSSIL for stochastic systems since, as previously mentioned, FOSSIL cannot handle stochastic systems. In fact, PRoTECT is the first tool in the literature to offer stochastic barrier certificates. Alternatively, the user can set a maximum degree and run computations in parallel up to this degree, returning the barrier certificate with the highest confidence. We call this the “Parallel” setting, as shown in Table 3, where $\lambda$ is once again fixed a-priori. We remind the reader that, unlike the deterministic setting, PRoTECT awaits barriers of all degrees to terminate, returning the one with the highest confidence. Consequently, stochastic case studies typically require notably longer completion times compared to the “One Shot” setting or parallelism for deterministic systems (discussed next), but at the gain of offering a higher confidence, e.g. example ex_nonlin1 has a confidence improvement of $32\%$. This is a trade-off the user should navigate based on their particular setting.
Remark 2.
In Table 2, the VDP case does not yield an optimized solution for $\lambda=1000$. However, by setting the minimum confidence level to $\phi=0.8$ (as shown in red-4 in Figure 2), feasible values for $\lambda$, $\gamma$, and $c$ can be identified with confidence exceeding the specified threshold. These results are omitted from Table 2 to maintain consistency in the comparison.
Finally, for completeness, we discuss the performance of PRoTECT running in parallel for the deterministic case studies, similar to the stochastic ones. PRoTECT simultaneously searches for a valid BC of varying degrees, such as 2, 4, and 6. Once a valid BC is found, the other threads are terminated. The parallelism introduces minimal overhead when computing multiple results, and we found that running the degrees in parallel is 19-27% faster than computing the same degrees (2, 4, 6) sequentially. The cases in Table 1 all have valid degree 2 barrier certificates, and we omit the parallelism results due to space constraints.